Item 3 Part A Bjorn is standing at z = 500 m . Firecracker 1 explodes at the origin and firecracker 2 explodes at z = 850 m The flashes from both explosions reach Bjorns eye at t At what time did Firecracker 1 explode? Express your answer using two significant figures. Submit My Answers Give Up Part B At what time did Firecracker 2 explode? Express your answer using two significant figures. Submit P Give Up
Answer #1


part(A)

time taken for flash at explosion1 to reach t = d/c = 500/(3*10^8) = 1.67 us

therefore the firecracker1 explode at time t1 = 4 - 1.67 us = 2.33 us <<<-------ANSWER


=================

part B

time taken for flash at explosion1 to reach t = d/c = 350/(3*10^8) = 1.17 us

therefore the firecracker2 explode at time t2 = 4 - 1.17 us = 2.83 us <<<-------ANSWER